### 17 Apr 2026 v1

# Co-Design of CNN Accelerators for TinyML using Approximate Matrix Decomposition

Abstract
The paradigm shift towards local and on-device inference under stringent resource constraints is represented by the tiny machine learning (TinyML) domain. The primary goal of TinyML is to integrate intelligence into tiny, low-cost devices under strict resource, energy, and latency constraints. However, the ultra-resource-constrained nature of these devices can lead to increased inference execution time, which can be detrimental in latency critical applications. At the same time, TinyML applications are often associated with sensitive data. As such, latency optimization approaches that rely on training samples are infeasible when such data is unavailable, proprietary, or sensitive, highlighting a pressing need for optimization approaches that do not require access to the training dataset and can be applied directly to pre-trained models. Replacing costly multiplications with more hardware-efficient operations, such as shifts and additions, has been proposed as an effective method for reducing inference latency. However, post-training power-of-two (Po2) approaches are scarce and, in many cases, lead to unacceptable accuracy loss.
In this work, we propose a framework that applies approximate matrix decomposition to a given CNN in order to optimize hardware implementations subject to strict constraints and without any need of re-training or fine-tuning steps. The genetic algorithm-driven framework explores different matrix decompositions and resulting multiplier-less CNN accelerator designs for FPGA targets. A comprehensive evaluation of different TinyML benchmarks demonstrates our framework’s efficacy in generating latency-optimized implementations that satisfy strict accuracy and resource constraints, achieving an average 33% latency improvement with an average accuracy loss of 1.3% compared to typical systolic array-based FPGA accelerators.
I Introduction
In recent years, the complexity of convolutional neural network (CNN) workloads has risen exponentially [28, 34], leading to a strong reliance on cloud-based processing. Although cloud offloading provides scalability, it significantly strains computational resources, resulting in increased energy consumption and raising concerns about privacy and costs. These challenges have driven a paradigm shift towards local, on-device, energy-efficient, and ultra-resource-constrained machine learning (ML) deployment, which is represented by the TinyML domain. TinyML embeds intelligence in far-edge, cost-effective applications such as wearable healthcare monitoring [16], always-on anomaly detection [25], and object detection [27]. Despite these advancements, TinyML systems are highly resource-constrained in both memory and compute, with microcontroller units (MCUs) or resource-constrained field-programmable gate arrays (FPGAs) widely employed as target platforms [29, 18, 5, 21, 22]. The inherent resource constraints of such devices can lead to increased inference latency, which can become critical in TinyML applications such as tiny unmanned aerial vehicles [19] or wearable biomedical devices [31], where inference must be performed under strict latency constraints. To enable deployment and reduce inference latency under such constraints, CNNs are designed to be shallower [3] and architecturally simpler [14], with a reduced number of parameters [13, 15]. Techniques such as pruning and quantization have been widely adopted to reduce the hardware footprint of ML models [12]. Similarly, TinyML acceleration has focused on developing hardware support for such techniques, including native computation at lower bitwidths [9] or sparse processing [30]. However, these approaches still heavily rely on multiplication operations, which naturally introduce performance and resource overheads.
Multiplier-less inference has emerged as a promising strategy, replacing resource-demanding multiplications with lightweight operations, such as shift-and-add. FPGAs are well-suited for this approach, offering design customizability and optimal support for such architectures. Our initial experimentation, on a rather small FPGA, shows that a variable signed right-shift operation requires $1.5\times$ fewer look-up tables (LUTs) than an equivalent multiplier, showcasing the potential resource savings of multiplier-less inference. Binary neural networks (BNNs) [37, 33, 32] and power of two (Po2)-based networks [35, 20, 24, 23, 8, 38] are notable examples, with the latter offering greater representation flexibility and potentially leading to higher achievable accuracy [35]. However, Po2 transformation often necessitates training in order to avoid any substantial accuracy degradation. As user privacy is paramount in TinyML scenarios, retraining may be infeasible due to sensitive, proprietary, or unavailable training data, which is typically the case in third-party or licensed models [1]. Rather, post-training approaches enable a rapid deployment and iteration without incurring the computational and energy costs of retraining, thereby reducing development overhead. On the other hand, current Po2 approaches that operate entirely in post-training scenarios [20, 11] are hardware-unaware, adhering to only accuracy constraints and neglecting the inherent resource constraints of TinyML. To the best of our knowledge, no existing work enables Po2-based CNN inference under strict resource and accuracy constraints without retraining.
In this work, we address these limitations by proposing the first post-training, hardware-aware co-design framework for Po2-based CNN inference on TinyML devices. Our framework leverages approximate weight matrix decomposition (WMD) to transform targeted CNNs into Po2 form without the need for retraining, jointly targeting minimal latency and accuracy loss. To that end, we design a programmable, multiplier-less CNN accelerator tailored to the decomposed CNN. The accelerator replaces traditional multiply-accumulate (MAC) units with lightweight shift-and-add processing elements (PEs) and is fully parametrizable based on our framework’s decomposition outputs. Our framework jointly explores Po2 transformation strategies and accelerator configurations. It uses our hardware-driven latency and resource analytical models to generate accuracy-latency Pareto-optimal solutions, subject to the resource constraints of TinyML. Evaluated on the established MLPerfTiny benchmark suite [3] and the Arty A7-100T FPGA board, our framework demonstrates an average latency improvement of 33% at just 1.3% average accuracy loss, compared to modern 8-bit accelerator designs.
Our novel contributions in this work are as follows:
  • We introduce the first post-training, hardware-aware co-design framework that jointly optimizes CNNs and Po2-based accelerator configurations for Pareto-optimal trade-offs between accuracy and latency.
  • We design a parametrizable, multiplier-less FPGA accelerator with shift-and-add PEs for efficient hardware deployment of CNNs.
  • Our approach demonstrates substantial latency gains at negligible accuracy loss compared to the state of the art, while adhering to the strict constraints of TinyML.
II Background and Related Work
In this section, we first provide the fundamentals for decomposing a given CNN weight matrix into an approximated power-of-two form in Section II-A, followed by a discussion of related work in Section II-B.
II-A Approximate Weight Matrix Decomposition
In [24] a mathematical weight decomposition technique is presented, which transforms CNN weights into Po2 form. The goal of the weight matrix decomposition (WMD) process is, given a weight matrix $W$, to obtain an approximate representation $\hat{W}$ where the multiplications have been replaced by shift-and-add operations. The first step of the process can be seen in Fig. 1a. The weight matrix of a CNN layer is first flattened in the kernel dimension $K^{2}$ to form the $N=K^{2}$ dimension, with the output channels $C_{out}$ concatenated to form $M$. After the $M\times N$-matrix $W$ is constructed, it is then split into a set of slices $\mathcal{S}$ with a slice width of $S_{W}$, such that $W=[W_{1}|W_{2}|...|W_{s}]$.
The decomposition process of each $W_{s}^{M\times S_{W}}$ matrix is illustrated in Fig. 1b. Each $W_{s}$ can be approximated using a set of $\mathcal{F}^{M\times M}$ matrices containing a total of $P$ decomposed matrices $F$, such that $W_{s}\approx F_{s,P}\ldots F_{s,1}F_{s,0}$. The initial matrix $F_{s,0}^{M\times S_{W}}$ has the same dimensions as $W_{s}$ and is composed of an identity matrix $I^{S_{W}\times S_{W}}$ and then padded with $(M-S_{W})$ zeroes. Then, each decomposed array $F$ is constrained to contain exactly $E$ non-zero elements, each with a range of $\pm 2^{Z}$ Po2 values, where $Z\in\mathbb{Z}$ represents the number of possible shifts. This WMD process can be fully described by the following parameters: $\{{P,Z,E,M,S_{W}}\}$, each one distinctly affecting the accuracy of the decomposed CNN.
Finally, Fig. 1c presents the decomposed version of a CNN convolution operation, assuming a convolutional kernel of size $N$ sliding over the input feature map. An input vector $In$ of size $N$ is also split into $S$ slices, such that $In=[In_{1}|In_{2}|...|In_{s}]$. A slice $In_{s}$ is then applied upon the decomposed matrices $F$ (corresponding to $W_{s}$) to produce $S$ partial sums. The convolution operation is then completed by accumulating the partial sums, as generated by the decomposed convolution, to produce $M$ outputs.
Figure 1: (a) CNN weight matrix transformation to $N$ and $M$ dimensions. (b) WMD of the transformed weight matrix to P decomposed F matrices with Po2 weights. (c) WMD multiplication with input and accumulation of the partial sums.
II-B Related Work
CNN inference acceleration has been widely studied, with numerous works proposing specialized compute engines, mostly based on 8-bit MAC units [6, 17]. Specifically for TinyML, ARM’s Ethos-U55 [2] is an embedded NPU targeting MCUs, supporting block sparsity and hardware support for a minimum precision of 8 bits. Similarly, [30] [30] integrated structured sparsity into vector execution on resource-constrained RISC-V cores to accelerate matrix multiplications. [26] [26] presented a streaming accelerator with layer-specific, mixed-precision PEs. [9] [9] provided native support for low-bitwidth execution of TinyML workloads on FPGAs. While these approaches optimize performance, they remain reliant on costly multipliers, which dominate resource utilization on constrained platforms.
Multiplier-less inference has been explored through BNNs [37, 33, 32] and Po2 quantization [35, 20, 24, 23, 8, 38], with Po2 methods offering more flexibility and potentially higher accuracy [35]. In [38], [38] applied quantization-aware training to transform CNN weights into Po2 values, while in DeepShift [8], [8] replaced multiplications with bitwise shifts and sign flips for both inference and training. [35] introduced ShiftAddNet [35], a model that separates shift and add layers to enhance accuracy. However, these techniques rely on retraining, which is infeasible when training data are proprietary, sensitive, or unavailable.
Several Po2-based approaches avoid retraining entirely. [23] [23] introduced WMD to approximate matrices using sparse and Po2 components. [20] [20] proposed a post-training Po2 quantization technique, used to implement an FPGA accelerator for small models with fully-unfolded datapaths and hardwired shifts and zeros. This was later extended to CNNs by [24] [24]. With ShiftCNN [11], [11] proposed a Po2 quantization algorithm alongside a Po2-based FPGA accelerator, while [36] [36] presented a multi-objective post-training reparametrization to minimize quantization error. However, these methods do not explicitly account for hardware resource constraints, often leading to designs that exceed the strict budget of TinyML devices and necessitate using large or even server-grade FPGA boards, making them impractical for far-edge deployment.
Our proposed framework addresses these gaps by jointly optimizing CNN weight decomposition and accelerator design to enable efficient, resource-compliant Po2 inference for TinyML devices without retraining. In Table I, we present a qualitative evaluation of our techniques, comparing them with state-of-the-art CNN inference approaches.
TABLE I: Qualitative taxonomy of the differentiating factors between our proposed approach and the state of the art.
Approach Multiplier-less Inference Without Retraining Hardware Aware
[2, 6, 17, 26]
[35, 8, 38]
[23, 20, 24, 11]
Ours
III Proposed Programmable Accelerator
We propose a programmable CNN accelerator for TinyML deployment on resource-constrained FPGAs. Our accelerator comprises a systolic array (SA), i.e., a 2-D grid of lightweight PEs, capable of supporting the WMD algorithm detailed in Section II-A and performing exclusively shift-and-add operations. Its programmability allows flexible mapping of computations, including folding workloads across multiple passes when necessary due to resource constraints. The details of the PE are described in Section III-A, and the SA arrangement and scheduling are presented in Section III-B.
III-A Processing Element (PE)
The microarchitecture of our PE is shown in Fig. 2a. Each PE contains pipelined sub-modules (referred to as F-blocks hereafter), computing the multiplication of an input vector $V_{in}$ by an F{F} matrix. In our PE, we implement $F_{0}$ and $F_{gen}$ as hard blocks, meaning that in the case of $P>2$, the $F_{gen}$ block is time-multiplexed. The PE computes the multiplication of one slice of the decomposed matrix as described in Section II-A, computing the equivalent of $S_{W}$ convolution operations per row, for a total of $M$ rows. Considering that each matrix F{F} is restricted to have exactly $E$ non-zero elements per row, each row of the F-block comprises $E$ shift units, which are then reduced using an adder tree. Moreover, the sparsity within the F{F} matrices is unstructured, meaning that each shift unit also requires a multiplexer to select which element of the $V_{in}$ is passed to the shift unit. The shift unit itself is comprised of $Z$ predefined shift values, which are selected using multiplexers.
Moreover, we take advantage of some observed properties of the F{F} matrices: First, ${F}_{s,1}$ matrices always contain their non-zero elements in their first $S_{W}$ columns, which allows for omitting position encoding bits in the signals and multiplexers before the multiplier inputs. Second, in an effort to reduce indexing cost, we replace one of the non-zero elements $E$ with a fixed value of $1$ on the diagonal of the $F$ matrix (referred to as diagonal optimization hereafter). The activation is then passed directly to the adder tree, which allows for encoding E1E-1 elements per row. Then, all F-blocks corresponding to these matrices just need E1E-1 multiplier units per row, while each element of the input vector is hardwired to the adder tree of the corresponding row. Finally, for the purpose of lowering hardware resources, our design considers a set of Po2 with only negative exponents, which allows implementing right-shifts only.
III-B Systolic Array (SA)
Figure 2: (a) Microarchitecture of our WMD PE. (b) Our systolic array (SA), composed of pipelined PEs, handling $PE_{x}\cdot PE_{y}\cdot S_{W}\cdot M$ convolution operations every cycle.
Our PEs are arranged in a 2-D grid in a pipelined manner, forming the core SA of our accelerator architecture, as shown in Fig. 2b. To map a convolution operation to our accelerator, we follow a process similar to Section II-A, but substituting $N$ for $C_{in}$. As can be seen in Fig. 2b, the input features reside in an on-chip input buffer, and then $S_{W}$ inputs flow from the top of the SA. Since each row corresponds to different output channels, the input features are shared along the y-axis. The partial sums are reduced along the x-axis via internal accumulators within each PE. They are then summed up with the outputs from the previous iteration and stored in an on-chip output buffer. The PEs in the x-axis are connected using $M$ adders. The parallelism of our SA (i.e., useful convolution MACs per cycle) can be determined based on the total number of PEs and the parallelism within each individual PE. At the PE level, this depends on the WMD parameters $S_{W}$ and $M$. Then the total number of outputs equivalent to MACs per cycle of our accelerator is $S_{W}\cdot M\cdot PE_{x}\cdot PE_{y}$, with $PE_{x}$ the number of columns and $PE_{y}$ the number of rows of the SA.
Figure 3: Our framework methodology. The input is a pre-trained CNN, the design space, and additional accuracy and latency constraints. Candidate solutions are evaluated in terms of accuracy and latency, resulting in a front of Pareto-optimal solutions.
Algorithm 1 PE Mapping
[…]
5:while $R_{PE_{x}}\leq LUT_{max}$ do
6:  $PE_{y}\leftarrow\left\lfloor{LUT_{max}}/{R_{PE_{x}}}\right\rfloor$
7:  $Map_{PE}\leftarrow(PE_{x},PE_{y})$
8:  $Lat\leftarrow Lat_{accl}(Map_{PE})$
9:  if $Lat<Lat_{best}$ then
10:   $Map_{best}\leftarrow Map_{PE}$
11:   $Lat_{best}\leftarrow Lat$
12:  end if
13:  $PE_{x}\leftarrow PE_{x}+1$
14:end while
15:return $Map_{best}$, $Lat_{best}$
To minimize data movement, we fetch the elements of all F{F} matrices from on-chip memory into registers within the PEs, and then traverse through the input features to be multiplied. This approach is analogous to the weight-stationary scheduling used in other SA-based CNN accelerators (e.g., TPU [17]). For the loading of the corresponding F{F} matrices, a total of $PE_{x}\cdot PE_{y}\cdot P\cdot E\cdot M$ elements need to be fetched and stored in registers. Our SA makes use of block random access memory (BRAM) resources available in FPGAs for the input and output buffers. One BRAM is used to feed each column and ${PE_{y}\cdot M\cdot out_{bw}}/{b_{ports}}$ BRAMs to store the packed results from all rows of the grid, with $out_{\textit{bw}}$ the bit-width of each output element and $b_{\textit{ports}}$ the summed bit-width of all ports of the BRAM.
To determine the 2D mapping $(PE_{x},PE_{y})$ on our SA, we use Algorithm 1. Starting from $PE_{x}=1$, the algorithm iteratively increments the length of the row, calculates $PE_{y}$ based on the total number of LUTs of the FPGA, and computes the latency $Lat$, based on the hardware cost of a single PE unit $R_{PE_{unit}}$ and the cost of a PE row $R_{PE_{x}}$, which can be derived from the F-blocks cost $R_{F_{gen}}$ and $R_{F_{0}}$ as described in Section IV-B. The process continues until the cost of one row exceeds the available LUTs. Once the algorithm terminates, the best latency $Lat_{best}$ is returned, ensuring that for a given CNN, the assignment best suited to its characteristics is chosen. This ensures that the best PE mapping is chosen for a given CNN based on the $C_{out},C_{in}$ dimensions.
IV Proposed Framework
In this section, we describe our novel hardware-aware framework for co-designing Po2-based CNNs and dedicated accelerators. Figure 3 presents an overview of our framework. It takes as input a pre-trained CNN, the design space of WMD parameters for each layer, and predefined application constraints for accuracy and latency. Then, candidate decomposed CNNs and accelerators are explored—aided by our fully parametrizable architecture (see Section III)—employing high-level surrogate models for latency and resource utilization for the targeted FPGA. Our framework outputs a Pareto front of Po2-decomposed CNN-accelerator configurations with optimal latency-accuracy trade-offs, without requiring any additional training.
IV-A Design Space
The WMD algorithm is fully parametrized, given the following variables: $P,Z,E,M,S_{W}$ (see Section II-A). Thus, a design space of different WMD-related choices can be defined, each affecting the accuracy, latency, and resources of the CNN accelerator. The set of WMD parameters is partitioned into a set of hard accelerator parameters $P_{h}$—applied to all decomposed layers LL—and a set of soft, layer-specific choices $P_{s}$ applied individually to each different layer, setting a design space of cardinality $|P_{s}^{L}|\times|P_{h}|$. We choose to partition the design space in this way, primarily because of its complexity. For a given CNN, the number of layers to be decomposed is constant; therefore, it has a polynomial dependency on $L$. Using this split, we can therefore have control over the complexity of the design space. Secondly, allowing each layer to be decomposed with $P_{s}$ WMD parameters means that the hardware needs to be able to support all choices made by each layer $L$, which can cause significant overhead in our programmable accelerator. In short, partitioning the design space in this manner allows us to control the cost in terms of hardware as well as the size of the design space. Although different partitions of WMD parameters can be considered, for the rest of this paper we set the number of decomposition blocks to be the only soft parameter ($P_{s}=\{P\}$) and the rest to be hard parameters $P_{h}=\{Z,E,M,S_{W}\}$.
IV-B Accelerator Modeling
IV-B1 Resource Model
The total resources $R_{accl}$ for our accelerator, in terms of FPGA LUTs, can be estimated using Eq. 1. This value is dependent on the resources of $F_{0}$ and $F_{gen}$, the total number of PEs $PE_{x}$, $PE_{y}$, as well as the reduction cost between the PEs in the x-dimension.
$R_{accl}=PE_{y}PE_{x}(R_{F_{0}}+{R_{F_{gen}}}+R_{add}M)$ (1)
The resources for an F-block are detailed in Eq. 2 and Eq. 3. We begin our resource analysis from the resources $R$ of the base units used within our PEs. Then, we estimate the resources of our PE by multiplying the base cost by the number of times each block is instantiated. As described in Section III-A for one F-block, we use a Po2 shift unit, a multiplexer, and an adder tree with base resources of $R_{mul}$, $R_{mux}$, which are dependent on the supported shifts $Z$ and the bit-width of the maximum supported WMD stage $F_{max}$, respectively. Similarly, $R_{add}$ is dependent on the number of $E$ elements to be reduced and $F_{max}$. Moreover, we also account for the simplifications of the first F-block $F_{0}$ and the diagonal optimization in $F_{gen}$. Note that our resource models use direct hardware measurements, obtained from EDA tools for the target FPGA, and thus provide highly accurate estimations guiding the optimization process.
$\begin{split}R_{F_{gen}}(E,F_{max},Z)=M(E-1)(R_{mul}+R_{mux})+R_{add})\end{split}$ (2)
$\begin{split}R_{F_{0}}(S_{W},F_{0},Z)=M(S_{W}R_{mul}+R_{add})\end{split}$ (3)
IV-B2 Latency Model
As described in Section III-B, our accelerator offers parallelism in two distinct ways with a total parallelism of $(S_{W}\cdot PE_{x},M\cdot PE_{y})$ in the x- and y-dimension, respectively. We estimate the latency of the accelerator based on the total operation count within each convolution. This is dependent on the size of the output $O_{x,y}$, the kernel size $K_{x,y}$, and the input and output channels $C_{in}$ and $C_{out}$. $C_{in}$ is parallelized along the x-dimension and $C_{out}$ along the y-dimension. The final latency of the accelerator can be derived by dividing $C_{in}$ and $C_{out}$ by the total parallelism in the x- and y-dimension, respectively. Additionally, as detailed in Section III-B, our accelerator has support for the $F_{0}$ and a generic block $F_{gen}$, meaning that for $P$ total decompositions, there is hardware support for implementing the first two blocks $F_{0},F_{gen}$ in one cycle, while the rest is time-multiplexed over $F_{gen}$. This results in an additional latency factor of $Lat_{F}=1+(F_{max}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2})$. The total accelerator latency can therefore be derived by summing the individual layer latencies.
$Lat_{accl}=\sum_{l=1}^{L}{Lat_{F}{K_{x,y}}{O_{x,y}}\left\lceil\frac{C_{in}}{S_{W}\cdot PE_{x}}\right\rceil}\left\lceil\frac{C_{out}}{M\cdot PE_{y}}\right\rceil$ (4)
IV-C Genetic-Based DSE
The core of our framework is a hardware-aware two-objective optimization process, aiming to balance CNN accuracy with inference latency in a search space that is additionally restricted by the strict constraints of a maximum accuracy drop $Ad_{max}$ and a minimal latency $Lat_{std}$ imposed by the TinyML environment. The optimization is formalized in Eq. 5, where we define our objective function as a joint minimization problem of the accuracy drop $F_{acc}$ of the decomposed CNN and the latency of the accelerator $F_{lat}$. To ensure the validity of each solution $x$, we also constrain our optimization based on a maximum accuracy drop of $Ad_{max}$ and a minimum latency $Lat_{std}$. For a given solution $x$:
[…]
In order to tackle the optimization problem defined in Eq. 5, we employ a genetic algorithm, specifically the popular non-dominated sorting genetic algorithm (NSGA-II) [7]. Evolutionary-based exploration enables the effective traversal of complex design spaces (as in our case, see Section IV-A) in a reasonable amount of time, as they can be massively parallelized, thereby expediting the overall execution. NSGA-II in particular, due to its elitist nature, ensures that the best performing (i.e., fittest) solutions are retained throughout the optimization process, and a front of Pareto-optimal accuracy-latency trade-offs can be obtained upon convergence.
Candidate solutions (i.e., chromosomes) represent discrete samples of our design space, and are encoded in integer values. Starting from a randomly initialized population, the NSGA-II flow involves a sequence of fitness evaluations, uniform crossover, and random mutation operations, which guide the evolution to higher-quality solutions. Fitness evaluation follows the formulation in (5). Note, that for our genetic search we use only a small subset of the test set and reserve the majority of the test set to evaluate the final accuracy of our framework solutions. For ensuring exact accuracy calculations, we execute inference upon the approximate and deconstructed convolutional layers. First, we apply WMD to the target CNN layers individually, according to the decoded chromosome parameters (i.e., $P_{h}$ and $P_{s}$ of Section IV-A). The resulting set of decomposed matrices $\mathcal{F}$ is then used to reconstruct the approximate convolutional layers and the Po2-based CNN, following the reverse procedure, and allowing for directly executing inference with the approximate weights. To determine $F_{lat}$, we consider our high-level latency surrogate model described in Section IV-B2 to accurately quantify the total cycles using Eq. 4. Importantly, latency estimation is preceded by the iterative mapping algorithm used to arrange our PEs into the systolic architecture. This step utilizes our resource model (see Section IV-B1) to maximize utilization and deliver the lowest possible latency allowed by the selected set of WMD parameters.
Overall, our genetic-based exploration enables resource-aware co-optimization of Po2-based CNNs with programmable accelerators, generating Pareto-optimal trade-offs between bounded inference accuracy and latency.
V Results and Analysis
This section shows the evaluation of our framework, reporting the solutions obtained for an exploration run with constraints taken from an 8-bit SA using multipliers. We describe the benchmark and tools used for our experiments in Section V-A, and the results obtained in Section V-B. To evaluate our programmable accelerator architecture with WMD PEs considering hardware performance metrics, we also compare the obtained solutions against post-training, hardware-optimization techniques: post-training quantization (PTQ) with varying weight bit resolutions in Section V-C and a state-of-the-art FPGA implementation of another Po2 approach in Section V-D.
V-A Experimental Setup
We evaluate our framework on the set of CNNs available in MLPerfTiny [3], a widely used TinyML benchmark that comprises representative CNN workloads customized to the stringent TinyML resource constraints. As CNN benchmarks, it includes ResNet, MobileNetV1, and DS-CNN, trained on CIFAR-10, VWW, and Speech Commands, respectively. For the framework evaluation, we exclusively use the pre-trained models of all three CNNs from MLPerfTiny, without additional retraining or finetuning. TFLite is used to evaluate the accuracy of both our decomposed and the baseline CNNs. We evaluate the solutions found with our framework against traditional state-of-the-art SA-based accelerators [17] that use low precision MAC units (i.e., weights from 4 to 8 bits) across all MLPerfTiny CNNs. The mapping of each CNN to these SAs is obtained using Algorithm 1. Each row and column of the SA can be supplied with weights and activations in a single clock cycle through appropriately instantiated BRAM-based input and output buffers. As our target platform, we use the Artix-7 XC7A100TCSG324-1 FPGA. A fully parametrized SA using our WMD PEs as well as the SAs using MAC units as PEs are implemented in SystemVerilog. Vivado 2023.1 is used for simulation and synthesis to obtain FPGA resource utilization in terms of LUTs, flip-flops (FFs), and BRAMs, as well as to estimate power. For our resource model, we synthesize its basic blocks and extract their resource utilization in terms of LUTs (see Section IV-B1). The final latency is calculated for each CNN on a layer-wise basis using Vivado simulation, from which toggle rates are also obtained to serve as inputs for the power consumption analysis. Functional verification uses uniformly random inputs, decomposed weights, and $8$-bit activations.
Our NSGA-II is implemented using pymoo [4], with an initial population of $250$ chromosomes, running for up to $20$ generations. We set the crossover probability and degree of mutation $\eta_{c}$ to $0.9$ and $15$, respectively, and the mutation $\eta_{m}$ to $5$. For the genetic exploration we reserve $10\%$ of the test set for the exploration itself and $90\%$ of the test set to evaluate the final accuracy of our designs. For our largest benchmark (i.e., MobileNet), the size of the explored design space—as defined by the WMD parameters $P_{h},P_{s}$—reaches for $P_{s}=54$, $2^{12}\cdot 54=221\,184$ possible solutions, highlighting the need for employing genetic-based exploration within our framework.
V-B Evaluation
Figure 4: Pareto front obtained from NSGA-II exploration for different network models showing accuracy loss and normalized speed-ups $S=Lat_{\mathrm{s}td}/Lat(x)$ for each found Pareto point $x$.
(a) PTQ results for DS-CNN
Refer to captionRefer to caption   Refer to captionRefer to caption
(b) PTQ results for ResNet
Refer to captionRefer to caption   Refer to captionRefer to caption
(c) PTQ results for MobileNet
Refer to captionRefer to caption   Refer to captionRefer to caption
Figure 5: Post-training quantization (PTQ) comparison for (a) DS-CNN, (b) ResNet, and (c) MobileNet. The graphs show the normalized GOPS on the left and the accuracy drop on the right.
TABLE II: Detailed comparison of DS-CNN from our proposed framework against systolic arrays of different weight resolution.
MAC-based SA Ours
Weight Resolution 4-bit 5-bit 6-bit 7-bit 8-bit 8-bit1
Top-1 Accuracy (%) 80.83 91.05 91.67 92.36 91.95 90.63
Number of LUTs (overall utilization) 62531 (99%) 62586 (99%) 62825 (99%) 61873 (99%) 61612 (97%) 59922 (95%)
Number of 36-Kb BRAMs (overall utilization) 18 (13%) 18.5 (14%) 37.5 (28%) 13.5 (10%) 7.5 (6%) 11 (8%)
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25900 (20%) 21233 (17%) 22927 (18%) 19033 (15%)
Frequency (MHz) 125 113 122 111 114 122
Latency ($\bm{\mu}$s) 21.02 22.99 31.34 32.93 30.79 16.88
Peak Throughput (GOPS) 156 116 109 81 89 187
Estimated Power (mW) 1166 1117 1067 1060 1097 1237
Estimated Energy ($\bm{\mu}$J) 24.51 25.68 33.43 34.91 33.77 20.88
Estimated Energy Efficiency (GOPS/W) 133.59 104.00 102.37 76.36 81.13 151.17
  • 1 PW-Conv(1–4), PP=2, ZZ=3, EE=3, MM=4, $S_{W}$=4
TABLE III: Detailed comparison of ResNet from our proposed framework against systolic arrays of different weight resolution.
MAC-based SA Ours
Weight Resolution 4-bit 5-bit 6-bit 7-bit 8-bit 8-bit2
Top-1 Accuracy (%) 53.55 75.31 85.16 86.47 87.08 85.63
Number of LUTs (overall utilization) 62531 (99%) 62586 (99%) 62415 (98%) 61873 (98%) 60757 (96%) 55450 (87%)
Number of 36-Kb BRAMs (overall utilization) 18 (13%) 18.5 (14%) 13 (10%) 13.5 (10%) 13.5 (10%) 54 (40%)
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25427 (20%) 21233 (17%) 22618 (18%) 13674 (11%)
Frequency (MHz) 125 113 123 111 113 114
Latency ($\bm{\mu}$s) 236.80 259.74 278.12 306.82 302.58 250.24
Peak Throughput (GOPS) 155.77 116.17 111.39 80.95 84.57 117
Estimated Power (mW) 1166 1117 1064 1060 1041 1709
Estimated Energy ($\bm{\mu}$J) 276.11 290.13 295.92 325.23 314.98 427.66
Estimated Energy Efficiency (GOPS/W) 133.59 104.00 104.69 76.36 81.24 68.46
  • 2 Conv3x3(1–7), PP=2, ZZ=3, EE=3, MM=16, $S_{W}$=4
TABLE IV: Detailed comparison of MobileNet from our proposed framework against systolic arrays of different weight resolution.
MAC-based SA Ours
Weight Resolution 4-bit 5-bit 6-bit 7-bit 8-bit 8-bit3
Top-1 Accuracy (%) 78.33 81.51 83.11 84.18 84.07 82.88
Number of LUTs (overall utilization) 62531 (99%) 62586 (99%) 62959 (99%) 61873 (98%) 62367 (98%) 62506 (99%)
Number of 36-Kb BRAMs (overall utilization) 18 (13%) 18.5 (14%) 15.5 (11%) 13.5 (10%) 15 (11%) 37.5 (27.7%)
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25545 (20%) 21233 (17%) 23170 (18%) 18390 (17.5%)
Frequency (MHz) 125 113 123 111 113 114
Latency ($\bm{\mu}$s) 100.34 115.05 124.59 149.65 147.99 87.20
Peak Throughput (GOPS) 155.77 116.17 112.37 80.95 86.83 161
Estimated Power (mW) 1166 1117 1086 1060 1073 $17$
Estimated Energy ($\bm{\mu}$J) 116.99 128.52 135.30 158.63 158.79 149.72
Estimated Energy Efficiency (GOPS/W) 133.59 104.00 103.47 76.36 80.92 93.77
  • 3 PW-Conv(2–13), PP=2, ZZ=3, EE=3, MM=8, $S_{W}$=4
First, we evaluate the effectiveness of our proposed framework in delivering Pareto-optimal solutions in terms of classification accuracy drop $F_{\mathrm{a}cc}(x)$ in percent using the accuracy of the corresponding floating-point trained network model, and normalized speedups $S=Lat_{\mathrm{s}td}/Lat(x)$ for each found Pareto point $x$ with $Lat_{\mathrm{s}td}$ corresponding to the latency of an 8-bit baseline SA. Figure 4 presents the obtained Pareto fronts from our NSGA-II-based exploration across all studied benchmarks.
Overall, we observe a wide range of trade-offs between latency and accuracy, covering both ends of the spectrum—from solutions that prioritize high accuracy to those achieving significant latency gains. Most of the found Pareto-optimal solutions fall within the region of a conservative 2% accuracy drop, reaching up to 34% estimated speed-up over the baseline. Interestingly, our framework discovered a decomposed Po2-based version of DS-CNN, which improves its accuracy by 1% during an iso-latency evaluation. This demonstrates the effectiveness of our framework in exploring potential decomposed CNNs and accelerator pairs, enabling Po2 inference with small accuracy loss and notable latency gains.
Next, we perform an in-depth analysis of selected solutions from each Pareto front of Fig. 4 in terms of accuracy, hardware utilization, latency, throughput, and power, for the targeted FPGA board. Specifically, we select the highest-performing decomposed CNN with an accuracy loss below the 2% threshold. Table II, Table III, and Table IV present the obtained results in comparison with the $n$-bit SAs across all benchmarks. While this section focuses on the 8-bit SA comparison against the WMD PEs accelerator solutions obtained with its latency constraint as input to our framework, a detailed comparison with all bit-width SAs variants is discussed in Section V-C.
[…]
Figure 6: Comparison of SA and obtained accelerators’ performance metrics showing latency, accuracy, and energy efficiency trade-offs across models.
V-C Comparison Against State-of-the-Art Low-Precision MAC-based Systolic Array
We compare our approach against PTQ, a state-of-the-practice technique for improving hardware efficiency [10]. Using the same baseline SA architecture and adjusting the precision of the MAC unit, we progressively quantize the weights of the evaluated CNNs to lower bit-widths while keeping activations fixed at $8$ bits. Figure 5 shows the achievable peak throughput of the MAC-based SA at various bit-widths, normalized with respect to the accelerators obtained with our framework. Overall, we observe that PTQ at bit-widths lower than $5$ bits incurs substantial accuracy loss, well over the $2\%$ accuracy drop threshold. At the same time, our approach consistently outperforms PTQ within the $2\%$ accuracy drop threshold showcasing the effectiveness of our Po2-based framework in providing throughput improvements at a minimal loss in terms of accuracy. Table II, Table III, and Table IV contain the PTQ obtained top-1 accuracies, hardware resources and performance metrics for the different weight resolution systolic arrays. The best values for each row are highlighted for an easier comparison. The 4-bit SA, although providing advantages in many hardware performance metrics, causes a top-1 accuracy drop of at least 6% for the analyzed CNNs, which is why this configuration is excluded in the following comparisons. It can be observed that the DS-CNN obtained configuration outperforms the SAs across most of the hardware performance metrics. In general, the slightly higher estimated power consumption of the obtained accelerators is compensated by the higher peak throughput, which leads to lower latencies and higher energy efficiency in terms of GOPS/W. Figure 6 shows that the accelerator solutions found with our framework lie in the Pareto front of all evaluated CNNs, demonstrating optimal trade-offs between latency, accuracy, and energy efficiency.
[…]
V-D Comparison Against State-of-the-Art Post-Training Po2 FPGA Implementation
Finally, we compare our framework methodology to the state-of-the-art Po2 acceleration technique, ShiftCNN, proposed by [11] [11]. We re-implement ShiftCNN targeting the same FPGA as in our work and by precisely following the design directives detailed in [11]. This approach approximates a normalized weight by adding $N$ numbers from a set of negative powers of 2, each selected by a $B$-bit index. Then, the hardware architecture includes a precomputed tensor storing right-shifted input activations by different amounts, for which just one is selected using encoded $N\cdot B$-bit weights as inputs to multiplexers. A parameter $C$ is also considered in the architecture to define the number of weight and activation pairs that the accelerator can receive as inputs per cycle, which at the same time defines the number of multiplexers instantiated. While not explicitly described, $N\cdot C$ multiplexers are needed to keep the multiplication result in a single cycle. All multiplexers’ outputs are accumulated with an adder tree.
Figure 7: Comparison between our framework and Shift-CNN: GOPS (left) and accuracy drop (right).
TABLE V: Throughput and accuracy drop obtained by different ShiftCNN variants for the MLPerfTiny selected CNNs
Accuracy Drop
NN $B$ LUTs Frequency (MHz) Instantiable Adder Trees OPs GOPS DS-CNN ResNet MobileNet
4 2 11791 101 5 640 64.49 0.43 0.39 1.86
3 3 13793 93 4 512 47.58 1.53 0.14 6.22
3 2 9516 108 6 768 82.57 7.71 2.74 30.8
[…]
VI Conclusions
Despite the substantial benefits of multiplier-less Po2 acceleration for the TinyML domain, its implementation remains challenging. Training-based approaches offer high accuracy but are impractical when the training dataset is unavailable. On the other hand, existing post-training approaches for achieving Po2 inference do not take into account the underlying hardware architecture, resulting in a high resource utilization.
To address these limitations, we proposed a novel post-training co-design methodology that accelerates tiny CNNs using Po2 representations via approximate WMD. Our approach explicitly incorporates hardware awareness, enabling flexible accuracy–latency trade-offs and achieving low-latency CNN inference with only minimal accuracy degradation. From evaluations using the PerfTinyML benchmark CNNs, we obtained an average of 33% latency improvement at an average accuracy loss of 1.3% compared to typical systolic-array-based FPGA accelerators. When compared with post-training-quantization hardware-optimized variants, the solutions obtained by our framework lie on the Pareto front of the accuracy drop and latency objectives used for the exploration. Additionally, a comparison with a state-of-the-art that also employs Po2 decomposition for CNN acceleration shows an average $2.4\times$ increase in achievable throughput with our hardware solutions, considering an accuracy drop of less than 2% across the analyzed CNNs.